# Scaling Analysis of Interleaved Speech-Text Language Models

### v1

@@ Abstract
Existing Speech Language Model (SLM) scaling analysis paints a bleak picture. They predict that SLMs require much more compute and data compared to text, leading some to question the feasibility of training high-quality SLMs. However, modern SLMs are often initialised from pre-trained TextLMs using speech-text interleaving to allow knowledge transfer. This raises the question - Do interleaved SLMs scale more efficiently than textless-SLMs? In this paper we answer a resounding yes! We conduct scaling analysis of interleaved SLMs by training several dozen and analysing the scaling trends. We see that under this setup SLMs scale more efficiently with compute. Additionally, our results indicate that the scaling-dynamics are significantly different than textless-SLMs, suggesting one should allocate notably more of the compute budget for increasing model size over training tokens. We also study the role of synthetic data and TextLM model families in unlocking this potential. Results suggest, that our scaled up model achieves comparable performance with leading models on speech semantic metrics while using less compute and data than other approaches. We open source models, samples, and data - https://pages.cs.huji.ac.il/adiyoss-lab/sims/.
@@ 1 Introduction
Recent methods for training SLMs integrate text alongside speech, either by processing both modalities concurrently (Défossez et al., 2024; Fang et al., 2024) or by interleaving between them (Nguyen et al., 2024; Zeng et al., 2024). These models are often referred to as joint speech-text SLMs. Such modelling approach resulted in significant improvement in SLM performance, specifically considering semantic content understanding and generation. Despite the great results, the scaling properties of such training paradigms remains unclear.
@@
In this work, we explore the scaling behaviour of joint speech-text SLMs within an interleaved setup. We train multiple interleaved SLMs across a range of sizes, compute budgets, and model families to derive practical insights for optimizing SLM performance. Specifically, after identifying well-performing model families and synthetic data mix - we fix several compute budgets and compare models of different sizes trained for the same compute.
@@
The insights found directly translate into better SLMs, which far out-perform the expected results when considering speech-only scaling laws (Cuervo & Marxer, 2024). This in turn gives an optimistic tone to the feasibility of training high-quality SLMs with existing compute and datasets. We apply all insights and train a 7B interleaved speech-text SLM which achieves performance comparable to leading SLMs considering all semantic speech metrics. We open source all the above models and training code, thus encouraging the community to further analyse such scaling properties, and text-speech knowledge transfer.
@@ 2.1 Speech Language Models
The speech research community has explored various types of SLMs (Cui et al., 2024b). In this work, we focus on models trained using both speech and text tokens—commonly referred to as joint speech-text SLMs. This approach typically involves three key components: (i) converting speech to discrete units, (ii) a joint speech-text SLM, and (iii) converting units back to speech, with each module trained independently. While previous studies introduced different methods for jointly modelling speech and text, we focus on modality interleaving.
@@
Different approaches to assigning modalities to time-spans or words have been suggested. We adopt a strategy similar to that of Zeng et al. (2024), where speech segments are chosen by sampling their lengths (in words) from a Poisson distribution with , continuing until these segments make up of the total word count. A language model (often initialised from TextLM) is then trained on the resulting mixed-modality sequences following the standard next-token prediction objective. Previous studies have also suggested training models on individual modalities - i.e. using separate speech-only and text-only datasets.
@@ 2.2 Scaling Analysis
where , , , , and are parameters empirically estimated using data from multiple trained models. The third approach varies the model size for a fixed set of different training FLOPs. It then identifies the optimal model size for each training FLOPs budget. Based on the optimal models, it fits a function to predict the optimal number of training tokens or model size given training FLOPs budget. This appraoch was termed ISO-FLOP curves.
@@ 3 Experimental Setup
[…]
@@
Notice, both sTinyStories, sSC, and tSC were synthesised using a single speaker dataset, specifically using LJ (Ito & Johnson, 2017) voice. This process is aligned with prior work (Hassid et al., 2023; Cuervo & Marxer, 2024). Hence, to assess how well models trained on single-speaker synthetic datasets generalise to other speakers, we construct a multi-speaker version of the spoken StoryCloze benchmark. Specifically, we use Kokoro text-to-speech (TTS) (Hexgrad, 2025) to synthesise the same texts from the original Spoken-StoryCloze and Topic-StoryCloze datasets with four distinct speakers. This setup also allows us to separately generate audio prefixes and suffixes, enabling cross-modal evaluation (e.g., audio prefix with a text suffix). We will release these benchmarks alongside the publication.
@@
4 Building the Scaffolding for Scaling
In order to perform a scaling analysis of SLMs (Section 5), we start by analysing the impact of the different design choices which could affect scaling experiments - model family, differences from non-interleaving and TWIST initialisation.
@@ 4.1 The Effect of Synthetic Data
Results provided in Table 1 indicate that models trained solely on synthetic data perform well on metrics involving the same single speaker (LJ)—such as sSC and tSC—but show weaker performance on evaluations involving other speakers, like sBLIMP and multi-speaker tSC. This is also evident from the validation cross-entropy on real, speech-only data. While prior research indicates that phonetic speech tokens carry limited speaker information (Sicherman & Adi, 2023), our results suggest that evaluating on a single speaker seen during training may give a misleading impression of model performance. Nevertheless, we see benefit across diverse speakers, by training on a mixture of synthetic and real data, compared to real data only or synthetic data only. This supports the choice of using a mixture of real and synthetic data. In the context of scaling laws, partially replacing real data with synthetic data reduces the speech data needed to train strong SLMs.
@@
We therefore suggest leveraging TTS-generated data from text corpora—providing more semantically coherent input—alongside real speech data to enhance SLM training. In the case of interleaved SLMs, it’s important to recognise that validation loss may not always reflect semantic performance accurately. Therefore, evaluation should include a range of metrics, particularly those involving varied and previously unseen speakers.
@@ 4.2 Model Family Impact
We conduct an empirical analysis leveraging leading TextLMs (B) as a base for interleaved SLMs. Specifically, we use LLama (Grattafiori et al., 2024), Qwen (Yang et al., 2024), Gemma (Team et al., 2024), SmolLM (Allal et al., 2025), Pythia (Biderman et al., 2023), OPT (Zhang et al., 2022), and Bloom (Workshop et al., 2022). Results, provided in Figure 2, suggest that not all model families are born equal. For instance, PythiaM underperforms OPTM despite its larger size. In addition, Qwen-B outperforms other, notably larger, models like SmolLM-B, Bloom-B and OPTB. We note that these results do not always correlate with performance of the base TextLMs. For instance SmolLM-B outperforms Qwen-B and LLaMa-B on standard text evaluations, yet, in our interleaved SLM setup, even the smaller Qwen–B model performs better.
@@ 4.3 When Does Text Benefit Speech Modelling?
Following the above observation that different model families (and sizes) perform very differently, we focus on a model which showed good performance - Qwen-B. We compare interleaved SLMs trained with the same recipe and initialisation, but varying number of steps. Our results, described in Figure 3, show that for semantic abilities the interleaved model outperformed the speech only model from as little as training steps ( tokens). This impact seems to increase with more compute. Other results show that this is not necessarily the case for weaker/smaller TextLMs, e.g. OPTM, which could require notably more steps for interleaving benefits, see Appendix 5.
Hassid et al. (2023) showed that initialising SLMs from a pre-trained TextLM could have diminishing returns with regards to training steps/tokens. Intuitively, the longer the model is trained, the less impact the good initialisation has, though not necessarily decaying to zero. We wish to investigate this in the context of interleaved SLMs. We hypothesise that this effect could have an impact on the functional form between and performance. While in existing SLM scaling analysis (Cuervo & Marxer, 2024) the loss in assumed to have an exponential relationship with , with TWIST initialisation this could decay faster, and not necessarily exponentially.
We investigate this by comparing an interleaved SLM initialised from scratch to the common text-initialised version. Results are depicted in Figure 3. We observe that the gap in loss does seem to decrease with more training steps, however the gap in tasks metric has less clear cut (though TWIST initialisation is clearly beneficial). This could mean that the impact of on model performance might decay faster, at least for some ranges and initialisations. This could induce a different scaling functional form, but intuitively this would mean that one should spend more of the compute on parameters222One could consider the compute used for training the TextLM, however, these models are widely available thus there is not likely a need to train them from scratch..
@@ 5 Scaling Analysis
Next, we analyse the relation between the downstream metric performance of the SLMs in relation to the training compute. We provide the results in Figure 1 and compare them to the estimates of textless-SLMs by Cuervo & Marxer (2024). We also fit an equation for the scaling of each, following the approach by Cuervo & Marxer (2024). However, we note that in our case these might be slightly more noisy, as we are limited with model sizes to existing TextLM sizes, our best model could be further from the actual optimal. Furthermore, certain scores could be near a saturation limit. Nevertheless, the results clearly show improved performance per-compute as well as a more optimistic trend. These enhance the feasibility of high semantic abilities in SLMs with existing data resources.
@@ 5.1 Comparison with existing models
We begin by evaluating the cross-modal capabilities of several of our models in Table 2, and compare their performance to SpiritLM, which operates with more than an order of magnitude more computational resources. Even though most of our analysis focused on speech only evaluation, we observe a similar trend for cross model evaluation tasks. Unsurprisingly, performance is the best for text-only. However, perhaps surprisingly the speech only performance sometimes out-performs cross-modal abilities. We hypothesise that the text modelling abilities are better than speech, yet-cross modality poses a challenge. Thus for complex enough tasks, e.g. sSC, the better text performance is enough to out-weigh the challenge of cross modality. However, for tSC this is not always the case.
@@ 6.1 Speech Language Models
The task of Generative Spoken Language Modelling over such speech representations was first introduced by Lakhotia et al. (2021). This was extended by Borsos et al. (2023), which incorporated both semantic and acoustic tokens to generate more expressive speech. Building on these foundations, many other SLMs have been proposed, each aiming to improve performance or address different aspects (Hassid et al., 2023; Baade et al., 2024; Maimon et al., 2025a; Cuervo & Marxer, 2024). Beyond the stream of SLMs we deal with, speech-aware TextLMs such as (Chu et al., 2023; Tang et al., 2023; Kong et al., 2024) have been proposed, where they train an audio encoder as an adapter to a pre-trained TextLM. SLMs were also used to solve various traditional speech tasks (Wang et al., 2023a; b; Popuri et al., 2022; Inaguma et al., 2022; Kreuk et al., 2021; Maimon & Adi, 2023; Roth et al., 2024). While many recent approaches incorporate speech-text in input and output (Défossez et al., 2024; Zhang et al., 2023; Cuervo et al., 2025; Zeng et al., 2024).
@@ 6.2 Scaling Analysis
Scaling analysis in TextLMs explores how performance improves with more computational resources, mainly model size and training data (Hestness et al., 2017; Clark et al., 2022; Kaplan et al., 2020). Hoffmann et al. (2022) introduced three scaling methods: varying training tokens with fixed model sizes, adjusting model size for fixed training FLOPs and fitting a parametric loss function (scaling law). Muennighoff et al. (2023) extended this work by examining how scaling laws apply to TextLMs in multi-epoch training scenarios. Additionally, Choshen et al. (2024b) analysed over 1,000 scaling laws, providing practical recommendations for conducting scaling research.
In the field of speech, several studies have explored scaling approaches, with a primary emphasis on scaling model size and data for multilingual automatic speech recognition (ASR) (Xiao et al., 2021; Pratap et al., 2020; 2024). Recently, Chen et al. (2025) investigated scaling laws in ASR, revealing that scaling enhances performance on low-resource languages. Most relevant to our work is Cuervo & Marxer (2024), which investigated scaling laws for SLMs in textless environment.
@@ Acknowledgements.
This research work was supported in by ISF grant number 2049/22, and by the NVIDIA Academic Grant Program providing computing resources, A100 and H100 GPUs hours on NVIDIA cloud computing cluster.
@@ A.1 Datasets
We use several speech datasets for training. Specifically, we use Libri-Light (Kahn et al., 2020), LibriSpeech (Panayotov et al., 2015), SWC (Baumann et al., 2018), Tedlium (Hernandez et al., 2018), PeopleSpeech (Galvez et al., 2021) and VoxPopuli (Wang et al., 2021). We only take English subsets for all datasets. We use full training splits where provided, otherwise splitting 99% for training. Full dataset sizes are in Table 4.

### v2

@@ Abstract
Existing Speech Language Model (SLM) scaling analysis paints a bleak picture. It predicts that SLMs require much more compute and data compared to text, leading some to question the feasibility of training high-quality SLMs. However, modern SLMs are often initialised from pre-trained TextLMs using speech-text interleaving to allow knowledge transfer. This raises the question - Do interleaved SLMs scale more efficiently than textless-SLMs? In this paper we answer a resounding yes! We conduct scaling analysis of interleaved SLMs by training several dozen and analysing the scaling trends. We see that under this setup SLMs scale more efficiently with compute. Additionally, our results indicate that the scaling dynamics significantly differ from textless-SLMs, suggesting one should allocate notably more of the compute budget to increasing model size over training tokens. We also study the role of synthetic data and TextLM model families in unlocking this potential. Results suggest that our scaled up model achieves comparable semantic speech performance to leading models, while using less compute and data. We open source models, samples, and data - https://pages.cs.huji.ac.il/adiyoss-lab/sims/.
@@ 1 Introduction
Recent methods for training SLMs integrate text alongside speech, either by processing both modalities concurrently (Défossez et al., 2024; Fang et al., 2024) or by interleaving between them (Nguyen et al., 2024; Zeng et al., 2024). These models are often referred to as joint speech-text SLMs. Such modelling approach resulted in significant improvement in SLM performance, specifically considering semantic content understanding and generation. Despite the great results, the scaling properties of such training paradigms remain unclear.
@@
In this work, we explore the scaling behaviour of joint speech-text SLMs within an interleaved setup. We train multiple interleaved SLMs across a range of sizes, compute budgets, and model families to derive practical insights for optimising SLM performance. Specifically, after identifying well-performing model families and data mix - we fix several compute budgets and compare models of different sizes trained with the same compute budget.
@@
The insights found directly translate into better SLMs, which far out-perform the expected results when considering speech-only scaling laws (Cuervo & Marxer, 2024). This in turn gives an optimistic tone to the feasibility of training high-quality SLMs with existing compute and datasets. We apply all insights and train a B interleaved speech-text LM which achieves performance comparable to leading SLMs considering semantic speech metrics. We open source the above models and training code, thus encouraging the community to further analyse such scaling properties, and text-speech knowledge transfer.
@@ 2.1 Speech Language Models
The speech research community has explored various types of SLMs (Cui et al., 2024b). In this work, we focus on models trained using both speech and text tokens, commonly referred to as joint speech-text SLMs. This approach typically involves three key components: (i) converting speech to discrete units, (ii) a joint speech-text SLM, and (iii) converting units back to speech, with each module trained independently. While previous studies introduced different methods for jointly modelling speech and text, we focus on modality interleaving.
@@
Different approaches to assigning modalities to time-spans or words have been suggested. We adopt a strategy similar to that of Zeng et al. (2024), where speech segments are chosen by sampling their lengths (in words) from a Poisson distribution with , continuing until these segments make up of the total word count. A language model (often initialised from TextLM; Hassid et al. 2023) is then trained on the resulting mixed-modality sequences following the standard next-token prediction objective. Previous works have also trained models on individual modalities - i.e. separate speech-only and text-only datasets.
@@ 2.2 Scaling Analysis
where , , , , and are parameters empirically estimated using data from multiple trained models. The third approach varies the model size for a fixed set of different training FLOPs. It then identifies the optimal model size for each training FLOPs budget. Based on the optimal models, it fits a function to predict the optimal number of training tokens or model size given training FLOPs budget. This approach was termed ISO-FLOP curves.
@@ 3 Experimental Setup
Optimisation. We follow the efficient SLM training recipe presented by Maimon et al. (2025a), and use their provided toolkit: SlamKit111https://github.com/slp-rl/slamkit. Specifically, we train a TWIST initialised SLM (Hassid et al., 2023), based on different TextLMs, using a cosine scheduler and the datasets described above (including the synthetic sTinyStories dataset). We provide full training hyper-parameters in Appendix A.2.
@@
Notice, both sTinyStories, sSC, and tSC were synthesised using a single speaker dataset, specifically using LJ (Ito & Johnson, 2017) voice. This process is aligned with prior work (Hassid et al., 2023; Cuervo & Marxer, 2024). Hence, to assess how well models trained on single-speaker synthetic datasets generalise to other speakers, we construct a multi-speaker version of the spoken StoryCloze benchmark. Specifically, we use Kokoro text-to-speech (TTS) (Hexgrad, 2025) to synthesise the same texts from the original Spoken-StoryCloze and Topic-StoryCloze datasets with four distinct speakers - male and female, British and American accents. Throughout this study we focus on the average performance across speakers, but provide per-speaker analysis in Appendix B.2.
This setup also allows us to separately generate audio prefixes and suffixes, enabling cross-modal evaluation, following the evaluation protocol of Nguyen et al. (2024); Cuervo et al. (2025); Zeng et al. (2024). Essentially, this is much like the uni-modal StoryCloze where we wish to see whether the ”real” continuation is assigned a higher likelihood by the SLM, compared to a distractor. However, for the cross-modal setting the prompt is one modality (e.g. text), and the continuation is a different modality (speech). In other words, given a specific text prompt the SLM needs to select the more likely speech continuation. We publicly release these benchmarks.
4 Building the Scaffolding for Scaling
To conduct a scaling analysis of SLMs (Section 5), we start by analysing the impact of key design choices which could affect scaling experiments - namely, the use of synthetic data, model family selection, and differences arising from interleaving and TWIST initialisation.
@@ 4.1 The Effect of Synthetic Data
Results provided in Table 1 indicate that models trained solely on synthetic data perform well on metrics involving the same single speaker (LJ)—such as sSC and tSC—but show weaker performance on evaluations involving other speakers, like sBLIMP and multi-speaker tSC. This is also evident from the validation cross-entropy on real, speech-only data. While prior research indicates that phonetic speech tokens carry limited speaker information (Sicherman & Adi, 2023), our results suggest that evaluating on a single speaker seen during training may give a misleading impression of model performance. Nevertheless, we see benefit across diverse speakers, by training on a mixture of synthetic and real data, compared to real data only or synthetic data only. This supports the choice of using a mixture of real and synthetic data. In the context of scaling laws, partially replacing real data with synthetic data reduces the real speech data collection needed to train strong SLMs. One could investigate synthetic data with diverse speakers, but we focus on the same single speaker as previous work and specifically Cuervo & Marxer (2024) for consistency.
@@
We therefore suggest leveraging TTS-generated data from text corpora, providing more semantically coherent input,alongside real speech data to enhance SLM training. In the case of interleaved SLMs, it’s important to recognise that validation loss may not always reflect semantic performance accurately. Therefore, evaluation should include a range of metrics, particularly those involving varied and previously unseen speakers.
@@ 4.2 Model Family Impact
We conduct an empirical analysis leveraging leading TextLMs (B) as a base for interleaved SLMs. Specifically, we use LLama (Grattafiori et al., 2024), Qwen (Yang et al., 2024), Gemma (Team et al., 2024), SmolLM (Allal et al., 2025), Pythia (Biderman et al., 2023), OPT (Zhang et al., 2022), and Bloom (Workshop et al., 2022). Results, provided in Figure 2, suggest that not all model families are born equal. For instance, PythiaM underperforms OPTM despite its larger size. In addition, Qwen-B outperforms other, notably larger, models like SmolLM-B, Bloom-B and OPTB. We note that these results do not always correlate with performance of the base TextLMs. For instance SmolLM-B outperforms Qwen-B and LLaMa-B on standard text evaluations (Allal et al., 2025), yet, in our interleaved SLM setup, even the smaller Qwen–B model performs better.
@@ 4.3 When Does Text Benefit Speech Modelling?
Following the above observation that different model families (and sizes) perform very differently, we focus on a model which showed good performance - Qwen-B. We compare interleaved SLMs trained with the same recipe and initialisation, but varying number of steps. Our results, described in Figure 3, show that for semantic abilities the interleaved model outperformed the speech only model from as little as training steps ( tokens). This impact seems to increase with more compute. Other results show that this is not necessarily the case for weaker/smaller TextLMs, e.g. OPTM, which could require notably more steps for interleaving benefits, see Appendix B.1.
Hassid et al. (2023) showed that initialising SLMs from a pre-trained TextLM could have diminishing returns with regards to training steps/tokens. Intuitively, the longer the model is trained, the less impact the good initialisation has, though not necessarily decaying to zero. We wish to investigate this in the context of interleaved SLMs. We hypothesise that this effect could have an impact on the functional form between and performance. While in existing SLM scaling analysis (Cuervo & Marxer, 2024) the loss in assumed to have an exponential relationship with , it is not clear this assumption holds in our setup. This training process resembles the addition of a new language to an existing LLM, hence its scaling behaviour might differ significantly. For example, the influence of tokens may not diminish exponentially due to factors like initialisation effects and potential knowledge transfer.
We investigate this by comparing an interleaved SLM initialised from scratch to the common text-initialised version. Results are depicted in Figure 3. We observe that the gap in loss does seem to decrease with more training steps, however the gap in tasks metric is less clear cut (though TWIST initialisation is clearly beneficial). This could mean that the impact of on model performance might decay faster, at least for some ranges and initialisations. This could induce a different scaling functional form, but intuitively this would mean that one should spend more of the compute on parameters.222One could consider the compute used for training the TextLM, however, these models are widely available thus there is not likely a need to train them from scratch.
@@ 5 Scaling Analysis
Next, we analyse the relation between the downstream metric performance of the SLMs in relation to the training compute. We provide the results in Figure 1 and compare them to the estimates of textless-SLMs by Cuervo & Marxer (2024). We also fit an equation for the scaling of each, following the approach by Cuervo & Marxer (2024). However, we note that in our case these might be slightly more noisy, as we are limited with model sizes to existing TextLM sizes, our best model could be further from the actual optimal. Nevertheless, the results clearly show improved performance per-compute as well as a more optimistic trend (the slope for sSC is higher, for tSC they are similar, but results are likely near the saturation limit). These results suggest the feasibility of high semantic abilities in SLMs with existing data resources.
@@ 5.1 Comparison with Existing Models
We begin by evaluating the cross-modal capabilities of several of our models in Table 2, and compare their performance to SpiritLM, which operates with more than an order of magnitude more computational resources. Even though most of our analysis focused on speech only evaluation, we observe a similar trend for cross-modal evaluation tasks. Unsurprisingly, performance is the best for text-only. However, perhaps surprisingly the speech only performance sometimes out-performs cross-modal abilities. We hypothesise that the text modelling abilities are better than speech, yet-cross modality poses a challenge. Thus for complex enough tasks, e.g. sSC, the better text performance is enough to out-weigh the challenge of cross modality. However, for tSC this is not always the case.
@@ 6.1 Speech Language Models
The task of Generative Spoken Language Modelling over such speech representations was first introduced by Lakhotia et al. (2021). This was extended by Borsos et al. (2023), which incorporated both semantic and acoustic tokens to generate more expressive speech. Building on these foundations, many other SLMs have been proposed, each aiming to improve performance or address different aspects (Hassid et al., 2023; Baade et al., 2024; Maimon et al., 2025a; Cuervo & Marxer, 2024). Beyond the stream of SLMs we deal with, speech-aware TextLMs such as (Chu et al., 2023; Tang et al., 2023; Kong et al., 2024) have been proposed, where they train an audio encoder as an adapter to a pre-trained TextLM. SLMs were also used to solve various traditional speech tasks (Wang et al., 2023a; b; Popuri et al., 2022; Inaguma et al., 2022; Kreuk et al., 2021; Maimon & Adi, 2023; Roth et al., 2024). Furthermore, many recent approaches incorporate speech-text in input and output (Défossez et al., 2024; Zhang et al., 2023; Cuervo et al., 2025; Zeng et al., 2024).
@@ 6.2 Scaling Analysis
Scaling analysis in TextLMs explores how performance improves with more computational resources, mainly model size and training data (Hestness et al., 2017; Clark et al., 2022; Kaplan et al., 2020). Hoffmann et al. (2022) introduced three scaling methods: varying training tokens with fixed model sizes, adjusting model size for fixed training FLOPs and fitting a parametric loss function (scaling law). Muennighoff et al. (2023) extended this work by examining how scaling laws apply to TextLMs in multi-epoch training scenarios. Additionally, Choshen et al. (2024b) analysed over scaling laws, providing practical recommendations for conducting scaling research.
In the field of speech, several studies explored scaling approaches, with a primary emphasis on scaling model size and data for multilingual speech recognition (ASR) (Pratap et al., 2020; Xiao et al., 2021; Pratap et al., 2024). Recently, Chen et al. (2025) investigated scaling laws in ASR, revealing that scaling enhances performance on low-resource languages. Most relevant to our work is Cuervo & Marxer (2024), which investigated scaling laws for textless SLMs.
@@ Acknowledgements.
This research work was supported in by ISF grant number , and by the NVIDIA Academic Grant Program providing computing resources, A and H GPUs hours on NVIDIA cloud computing cluster.
@@ A.1 Datasets
We use several speech datasets for training. Specifically, we use Libri-Light (Kahn et al., 2020), LibriSpeech (Panayotov et al., 2015), SWC (Baumann et al., 2018), Tedlium (Hernandez et al., 2018), PeopleSpeech (Galvez et al., 2021) and VoxPopuli (Wang et al., 2021). We only take English subsets for all datasets. We use full training splits where provided, otherwise splitting for training. Full dataset sizes are in Table 4.